OpenStudy (anonymous):

Match the power with its value. Match Term Definition eight to the second power A) 64 eight to the negative second power B) -64 negative eight to the second power C) -1/64 negative eight to the negative power D) 1/64 --------------------------------------------------------------------------------

OpenStudy (phi):

do you know what eight raised to the *positive* second power (written as \(8^2\) ) is ?

OpenStudy (anonymous):

no I was never good in math :/ I struggled a lot

OpenStudy (phi):

the little number in the top right means how many times to multiply itself for example \( 8^4\) means \( 8 \cdot 8\cdot 8\cdot 8\) , 8 times itself 4 times so what is \( 8^2 \) ?

OpenStudy (phi):

can you count to 2 ? when you see 8^2, write down 2 eights, and put a multiply sign in between can you do that ?